# TRMAC: A Time-Reversal-based MAC Protocol for Wireless Networks within Computing Packages

### v1

@@ 4 Methodology
The protocol leverages the unique characteristics of the on-chip wireless channel by starting from the physical layer. The framework of the methodology is illustrated in Figure 6. First, the chip package is designed and simulated using CST Microwave Studio [32] to extract the CIRs. These CIRs are pre-characterized and considered static, assuming a fixed chip layout and material composition. Second, MATLAB [33] simulations are used to implement a basic Amplitude Shift Keying (ASK) modulation and to calculate energy thresholds for bit detection with low BER. We then estimate the NPT that can coexist without exceeding acceptable BER limits, based on cross-interference among simultaneous links. Importantly, as discussed in Section 2.1, both the energy thresholds and NPT are fixed once the chip is designed and remain stable unless the physical environment changes. Finally, these parameters are incorporated into Multi2Sim [34] to evaluate protocol feasibility under practical operating conditions.
Figure 7: BER vs. data rate for parallel spatial transmissions using TR.
4.1 Physical Layer Simulations
@@
To model the on-chip wireless environment, a four-chiplet interposer package was simulated using CST Microwave Studio [32], with each core equipped with a single antenna operating at a center frequency of 140 GHz. The details of the flip-chip layers follows the setup described in [35]. The interposer is created by stacking four flip-chip structures on top of silicon interposer and copper layers. The vertical monopole antennas are placed on each chip-let in symmetrical positions such that the cylindrical metallic structure is passed through the silicon substrate and excited from the bottom with metal layer acting as the ground plane. After accurate antenna modeling, the CIR is obtained by sequential impulse excitations followed by the antennas, where we can extract the information of each link for post-processing. In MATLAB, serialized data is modulated using ASK, passed through a TR filter (complex conjugate of the CIR), and transmitted through the wireless channel. The receiver captures the signal and decodes the data by comparing received energy to pre-characterized energy thresholds [10, 11].
We extend this model to simulate multiple simultaneous transmissions (a MIMO scenario) to evaluate how BER varies with increasing transmission concurrency. For each transmitter, cross-interference from other links is analyzed. We determine the maximum number of spatial links (i.e. NPT) that can be reliably supported while maintaining BER below an acceptable threshold. In addition to interference, we analyze the influence of data rate. Higher data rates lead to tighter symbol spacing, increasing vulnerability to ISI, particularly in channels with high delay spread. The resulting impact of BER on date rate for multiple spatial links is presented in Figure 7. This highlights a critical trade-off between throughput and error in TR-based communication.
4.2 Wireless Transfer and Channel Allocation
Wireless data transfer in TRMAC can operate on a single or multiple frequency channels. This work focuses on a single frequency channel using three spatial channels enabled by TR. All transmissions are aligned to a global clock at 1 GHz.
Each packet consists of 80 bits: 15 for the preamble, 64 for data, and 1 for the tone as shown in Figure 5. This design supports addressing up to $2^{15}$ nodes. Communication is divided into 5 clock cycles: 1 for the preamble, 1 for ACK, and 3 for data, yielding a throughput of $80/5$ ns = 16 Gbps, which is well within the limits of achievable TR performance ($\sim$30 Gbps) with minimal CCI [10, 11].
The 140 GHz band (16 GHz bandwidth) is used for data transmission, while the 170 GHz band (1 GHz) carries the tone signal, as shown in Figure 8. This adds approximately 2 mW of power consumption to support a short bandwidth. The authors in [36], have implemented a OOK modulated non-coherant receiver obtaining 16Gb/s data rate at 60GHz antenna frequency, in 65nm CMOS. However, as TR filters have been explored from 60–200GHz center frequency [10, 12], the area and energy overhead will be approximately 1–5 mm2 and 2–5 pJ/bit/cm1/2 as evaluated in [37].
@@ 5 Evaluation
Number of antennas 64 – 256 (one antenna per chip)
@@
Network configuration 80 bits per packet, 15-bit preamble
@@
Standard deviation $\sigma=0.1$ – 10
@@
Center frequency 140 GHz (data), 170 GHz (tone signaling)
Bandwidth 16 GHz (data), 1 GHz (tone)
Data rate 30 Gbps (TR with ASK modulation)
@@ 5.1 Number of Parallel Transmissions
Figure 9 shows that TRMAC achieves comparable latency to BRS, and saturation throughput equivalent to Token Passing, despite operating on a single frequency channel. As the number of transmissions increases, latency in both TRMAC and BRS grows similarly due to increased contention. Token Passing exhibits the highest latency, which scales significantly with the ring length and added contention per token cycle.
@@
Figure 9: Latency-throughput curves for TRMAC and baselines with 64 cores using (a) two channels / NPT=2, (b) three channels / NPT=3 (c) four channels / NPT=4.
5.2 Number of Cores
We now evaluate the effect of increasing the number of cores (from 64 to 256) on protocol performance. We configure BRS and Token Passing with two frequency channels, while TRMAC uses NPT=4 on a single channel.
Figure 10 shows that as the number of cores increases while keeping per-core load fixed, total traffic increases proportionally. This leads to more collisions and an expansion of the token ring, degrading latency in Token Passing. Both TRMAC and BRS show lower latency at low injection rates, but Token Passing maintains higher throughput at high loads due to better collision avoidance.
However, in BRS and TRMAC, throughput slightly decreases with more cores due to increased contention. Despite this, TRMAC performs better than BRS at high injection rates, offering a favorable trade-off between latency and throughput as system size scales.
@@ 5.3 Traffic Sensitivity Analysis
We examine the sensitivity of TRMAC to spatial and temporal traffic variations using the synthetic traffic model. We analyze the standard deviation $\sigma$ (representing spatial distribution) and the Hurst exponent $H$ (representing temporal burstiness) across broad ranges: $\sigma=0.1$ 10 and $H=0.5$ 1. Results are compared with BRS, as both are designed for on-chip environments and exploit preamble-based access.
5.3.1 Spatial Distribution ($\sigma$)
@@
In Figure 11(a), TRMAC demonstrates reduced latency and increased throughput as NPT increases. More spatial channels allow better traffic dispersion. However, at higher $\sigma$, throughput slightly drops due to more concurrent access attempts causing contention.
Figure 11(b) shows that TRMAC retains stable performance as the number of cores increases from 64 to 256 under varying $\sigma$. Throughput and latency trends remain consistent, confirming robustness under spatially imbalanced traffic.
Figure 11(c) compares TRMAC with BRS across the full $\sigma$ range. Both protocols perform similarly under varying distribution, confirming TRMAC’s ability to handle both hotspot and evenly distributed traffic with efficient spectrum allocation.
Figure 11: Impact of spatial traffic distribution ($\sigma$) on (a) TRMAC 1 frequency channel with varying NPT, (b) TRMAC with increasing cores, one channel; NPT=3 (c) Comparison of TRMAC one frequency channel; NPT=3 and BRS three channels.
5.3.2 Temporal Burstiness ($H$)
@@
In Figure 12(a), TRMAC maintains similar throughput and latency for both NPT=2 and NPT=3, regardless of $H$. Figure 12(b) shows that even as core count increases, the performance of TRMAC remains stable across different burstiness levels.
Figure 12(c) compares TRMAC with BRS under bursty traffic. Both protocols show nearly identical performance trends, validating TRMAC’s robustness against temporal injection variability.
Figure 12: Impact of traffic burstiness ($H$) on (a) TRMAC with one channel and varying NPT, (b) TRMAC with increasing cores, one channel; NPT=3 (c) Comparison of TRMAC one channel; NPT=3 and BRS three channels.
@@ 6.2 Multiplexing
Multiplexing strategies have long been a foundational approach in WNoC MAC designs. Protocols based on TDMA, FDMA, and CDMA have been proposed to ensure low-contention data transmission through orthogonal resource allocation.

### v4

@@ 4 Methodology
The protocol leverages the unique characteristics of the on-chip wireless channel by starting from the physical layer. Therefore, the methodology of this work considers channel and physical layer simulations prior to the link-level simulations, as illustrated in Figure 7.
First, the chip package is simulated using CST Microwave Studio [34] to extract the channel responses. These channel responses are pre-characterized and considered static, assuming a fixed chip layout and material composition. On an actual implementation of the system, only a single-time calibration or measurement phase would be needed to obtain the channel for all antenna pairs. The channel is time-invariant, therefore those channel measurements are used during the lifetime of the device and there is no need to estimate the channel with repetitive probing.
@@
Second, MATLAB [35] simulations are used to implement a basic Amplitude Shift Keying (ASK) modulation and to calculate energy thresholds for bit detection with low BER. We then estimate the Number of Parallel Transmissions (NPT) that can coexist without exceeding acceptable BER limits, based on cross-interference among simultaneous links. Importantly, as discussed in Section 2.1, both the energy thresholds and NPT are fixed once the chip is designed and remain stable due to the static nature of the in-package channels. Finally, these parameters are incorporated into Multi2Sim [36] to evaluate protocol feasibility under practical operating conditions.
[…]
Figure 9: Total received power from A to the intended receiver B and non-intended receivers C and D
4.1 Physical Layer Simulation
@@
To model the on-chip wireless environment, a four-chiplet interposer package was simulated using CST Microwave Studio [34], with each core equipped with a single antenna operating at a center frequency of 140 GHz. The details of the flip-chip layers follows the setup described in [37]. The interposer is created by stacking four flip-chip structures on top of silicon interposer and copper layers. Vertical monopole antennas are placed on each chiplet in symmetrical positions such that the cylindrical metallic structure is passed through the silicon substrate and excited from the bottom with metal layer acting as the ground plane. The choice of a vertical monopole is motivated by its increased co-planar coupling. The length of the monopole is obtained by considering the antenna center frequency and the thickness and relative permittivity of the silicon substrate  [37]. The model we built consists of four 5×5 $mm^{2}$ chiplets as shown in Figure 8. The distances between the antennas are in the range of $1.29\lambda-4.38\lambda.$
After accurate antenna modeling, the CIR is obtained by sequential impulse excitations followed by the antennas, where we can extract the information of each link for post-processing. Theoretically, TR is compatible with any antenna design, as through the channel characterization entire channel response is obtained, including the antennas themselves, to calculate the filter’s response. In MATLAB, serialized data is modulated using ASK, passed through a TR filter (complex conjugate of the CIR), and transmitted through the wireless channel. The receiver captures the signal and decodes the data by comparing received energy to pre-characterized energy thresholds [10, 11].
With this model we simulate and analyze the cross-interference per each link by following Figure 2(a). The results in Figure 9 show us the received energy per each node in a single transmission with TR. When the signal is focused to the intended receiver as in link $AB$, the received power is high due to focused energy. However, a residual CCI is obtained as shown in links $AC$ and $AD$ and due to the cross-correlation between the links ABACAB-AC and ABADAB-AD. Yet, when multiple transmissions are utilized, the impact of CCI on each node is accumulated and consequently the BER is increased. However, with pre-characterized channel CIR we can determine the number of parallel transmissions (i.e. NPT) that can be reliably supported while maintaining BER below an acceptable threshold.
4.2 Link Layer Simulation
To assess the performance of TRMAC, several link-layer aspects need to be clarified and incorporated in Multi2sim. For instance, even though wireless data transfer in TRMAC can be performed over multiple frequency channels, this work focuses on a single frequency channel. The timing of all transmissions is considered to be aligned to a global system clock running at 1 GHz. Therefore, Multi2sim simulations are slotted at a nanosecond granularity. In this context, we consider the overall transmission to take $N+2$ clock cycles: one for the preamble transmission, one for the acknowledgment, and $N$ cycles for the transmission of the data.
In our particular set of simulations, each packet consists of 80 bits, which is consistent with the size of short control messages in multicore processors. Out of those 80 bits, 20 are for the preamble (address and first bits of data) and 60 bits for the rest of the data. Without loss of generality, we assume a wireless datarate of 20 Gbps so that the preamble fits within a single clock cycle. The rest of data can be transmitted in $N=3$ more clock cycles after the ACK. It must be noted that this datarate requirement is well within the limits of achievable TR performance ($\sim$30 Gbps) with minimal CCI that has been shown in prior work [10, 11].
We modified Multi2sim to integrate the TRMAC following the timing considerations above, and also taking into account that the tone signal prevents a node from starting the Phase I of a transmission during the Phase III of other nodes’ transmissions. To sustain both the data and tone signaling requirements, we use the 140 GHz band with 20 GHz bandwidth for data transmission and the 170 GHz band for the tone signal, as shown in Figure 10. We further note that the addition of the tone signal leads to a very small energy overhead. According to prior work, a tone signal at the chip level can add approximately 2 mW of power consumption. This results in a 0.075 pJ/bit of efficiency considering the three nanoseconds during which it must be active for each 80 bits of data transfer, which is negligible compared to the 1-pJ/bit target efficiency of WNoC.
@@ 5 Evaluation
Number of antennas 64–256 (one antenna per core)
@@
Network configuration 80 bits per packet, 20-bit preamble
@@
Spatial hotspotness $\sigma$ = 0.1–10
@@
Center frequency 140 GHz (data), 170 GHz (tone)
Bandwidth 20 GHz (data), 1 GHz (tone)
Data rate 20 Gbps (TR with ASK modulation)
Figure 11: Latency-throughput curves for TRMAC and baselines with 64 cores using (a) one channel for BRS and TOKEN / NPT=3 for TRMAC (b) two channels for BRS and TOKEN / NPT=2 for TRMAC, (c) three channels for BRS and TOKEN / NPT=3 for TRMAC (d) four channels for BRS and TOKEN / NPT=4 for TRMAC.
@@ 5.1 Number of Parallel Transmissions
We first observe the performance of TRMAC:3 NPT, with respect to single channel token/BRS. It is shown that in Figure 11(a), with similar channel conditions the latency of TRMAC is low and throughput is high with a high performance gain in compared to the baselines.
Consequently, we evaluate single-band TRMAC with multi-band token and BRS. Figure 11(b), Figure 11(c) and Figure 11(d) show that TRMAC achieves comparable latency to BRS, and much better latency than Token passing, despite operating on a single frequency channel. In terms of throughput, TRMAC achieves saturation throughput equivalent to token passing and equal performance as BRS with less resource allocation. As the number of transmissions increases, latency in both TRMAC and BRS grows similarly due to increased contention. Token Passing exhibits the highest latency, which scales significantly with the ring length and added contention per token cycle.
@@
5.2 Number of Cores
We now evaluate the effect of increasing the number of cores (from 64 to 256) on protocol performance. We configure BRS and Token Passing with 4 frequency channels, while TRMAC uses NPT=4 on a single channel.
Figure 12 shows that as the number of cores increases, total traffic increases proportionally. This leads to more collisions and an expansion of the token ring, degrading latency in Token Passing. Both TRMAC and BRS show lower latency at low injection rates, but Token Passing maintains higher throughput at high loads due to better collision avoidance.
However, in BRS and TRMAC, throughput slightly decreases with more cores due to increased contention. Despite this, TRMAC performs better than BRS at high injection rates with a single frequency channel while BRS, TOKEN uses 4 channels, offering a favorable trade-off between latency and throughput as system size scales.
Figure 13: Impact of spatial traffic distribution ($\sigma$) on the throughput (left) and latency (right) of TRMAC. Top plots compare the performance of TRMAC with one frequency channel and NPT=3 to that of BRS with three frequency channels. Bottom plots compare the performance of TRMAC with NPT=2 to that of TRMAC with NPT=3.
Figure 14: Impact of spatial traffic distribution ($\sigma$) on the throughput (left) and latency (right) of TRMAC when increasing number of cores with one frequency channel and NPT=3.
@@ 5.3 Traffic Sensitivity Analysis
We examine the sensitivity of TRMAC to spatial and temporal traffic variations using the synthetic traffic model. We analyze the standard deviation $\sigma$ (representing spatial distribution) and the Hurst exponent $H$ (representing temporal burstiness) across broad ranges: $\sigma=0.1$ - 10 and $H=0.5$ - 1. Results are compared with BRS, as both are designed for on-chip environments and exploit preamble-based access. The box plot representations follow a data set with the packet injection rates from 0.00001-1. The box itself represents the middle 50% data from first to the third quartile, while the line inside the box shows the middle value of the data set. The outside whiskers show the largest and smallest values within the range of 1.5 times the inter-quartile range.
5.3.1 Spatial Distribution
@@
The top plots of Figure 13 compares TRMAC with BRS across the full $\sigma$ range. Both protocols perform similarly under varying distribution, confirming TRMAC’s ability to handle both hotspot and evenly distributed traffic with efficient spectrum allocation such that having single frequency channel and NPT=2, 3 multiple parallel transmissions.
In Figure 13, TRMAC demonstrates increased throughput (bottom left) and reduced latency (bottom right) as NPT increases. More spatial channels allow better traffic dispersion. However, at higher $\sigma$, throughput slightly drops due to more concurrent access attempts causing contention.
Figure 14 shows that TRMAC retains stable performance as the number of cores increases from 64 to 256 under varying $\sigma$. Throughput and latency trends remain consistent, confirming robustness under spatially imbalanced traffic.
5.3.2 Temporal Burstiness
@@
Figure 15 (top plots) compares TRMAC with BRS under bursty traffic. Both protocols show nearly identical throughput trends, while the latency (top right) of TRMAC is smaller than BRS. This performance validates robustness of TRMAC against temporal injection variability and proving that it performs similarly with spatial diversity with one channel by using 1/2, 1/3 of resources compared to baseline protocols.
In Figure 15 (bottom plots), TRMAC maintains similar throughput and latency trend for NPT=2 and NPT=3, regardless of $H$. Figure 16 shows that even as core count increases, the performance of TRMAC remains stable across different burstiness levels.
Figure 15: Impact of spatial traffic burstiness ($H$) on the throughput (left) and latency (right) of TRMAC. Top plots compare the performance of TRMAC with one frequency channel and NPT=3 to that of BRS with three frequency channels. Bottom plots compare the performance of TRMAC with NPT=2 to that of TRMAC with NPT=3.
Figure 16: Impact of traffic burstiness ($H$) on the throughput (left) and latency (right) of TRMAC when increasing number of cores with one frequency channel and NPT=3.
@@ 6 Related Work
In this section, we review existing WNoC protocols, focusing on their medium access techniques, scalability, architectural design, and adaptability to varying traffic patterns. Table 2 summarizes several closely related and recent MAC protocols for WNoC, offering a comparative perspective on their performance characteristics.
@@ 6.1 Token Passing and CSMA
Table 2: Summary of the design traits of existing protocols
Work Architecture Multiplexing Scalability Traffic Awareness Remarks
Shamim et al. [4] Dynamic N/A High Low Energy efficient wireless links with token passing.
Mansoor et al. [19] Dynamic TDMA Medium High Time-slot allocation based on traffic prediction.
Franques et al. [20] Dynamic N/A High High Dynamic traffic control with token passing.
Mestres et al. [21] Dynamic N/A High Low Simple and highly adaptable for on-chip/package scenario.
Vijayakumaran et al. [13] Static CDMA High High Analytical evaluation of data transfer for on chip networks with CDMA, which leads to complex transceiver design.
Matolak et al. [14] Static FDMA, TDMA High Medium Hybrid wired-wireless architecture with low power and increased area efficiency.
DiTomaso et al. [17] Static TDMA High High Adaptable hybrid wired-wireless architecture improving energy efficiency
Jog et al. [23] Dynamic TDMA High High Deep reinforcement learning (DRL) based dynamic traffic control and a DRL framework to operate within precise timing, area and power constraints for on-chip communication networks.
Ollé et al. [22] Dynamic FDMA High High Exploring multi-channel medium access with existing protocols with dynamic traffic, random access and scalability.
Rout et al. [24] Dynamic N/A High High Congestion aware MAC with dynamic channel allocation according to the traffic requirements.
TRMAC (this work) Dynamic Spatial High High Time reversed parallel transmissions with spatial diversity in same frequency and time.
@@ 6.2 Multiplexing
Multiplexing strategies have long been a foundational approach in WNoC MAC designs. Protocols based on time, frequency, or code multiplexing (TDMA, FDMA, and CDMA) have been proposed to ensure low-contention data transmission through orthogonal resource allocation.
@@ 6.3 Architecture
7 Discussion and Limitations
TRMAC is a novel protocol which is designed for on-chip wireless networks that exploits TR filters for mitigating the channel impairments and to utilize parallel transmissions both in time and space. Despite of the spatial diversity, as TR filters do not prevent the transmission collisions, TRMAC fills the gap of coordination and control of TR precoded data transmissions. Unlike conventional MAC protocols that rely on frequency division or centralized arbitration, TRMAC operates using a single frequency channel while supporting multiple concurrent transmissions, thereby reducing both hardware complexity and spectral overhead.
The protocol prevents the collisions and addresses the deafness problem by transmitter-side channel sensing, acknowledgment feedback and tone-signaling in each phase of the proposed solution, by considering all the possible collision events that could occur with multiple concurrent transmitters and receivers operating in same time and frequency. Through comprehensive set of simulations it is shown that TRMAC maintains high throughput and low latency with multiple spatial channels, scales efficiently with increasing number of cores, adapt robustly to spatial and temporal traffic variations, while performing similar to or better than baseline protocols such as BRS and token passing with efficient resource allocation.
However, channel symmetry plays an important role in TRMAC, where the CIR of each link is composed with similar characteristics. Even though it is unlikely to obtain perfect symmetric channels with equal amplitude and phase, there is a possibility where the CCI could be increased due to spatial similarities. Yet, with the proposed TRMAC protocol, we pre-characterize each CIR prior to the transmission and set energy thresholds and NPT based on the evaluated CIRs. Channel symmetry could potentially limit the NPT, and the trade-off will be in between the expected NPT and the BER of the collective communications.
In terms of implementation, TR filters have to resemble a precise time-reversed CIR of the characterised channel response to mimic the ideal filter and to achieve perfect energy focusing. However, with the limitation of achievable sampling rates (up to a few tens of GSps), imperfect TR filters may negatively impact on the performance of the system. Additional limitations may come from the amount of memory required to store the filters. In extremely reverberant environments, the length of the channel could be several tens of nanoseconds long. Fortunately, even assuming very high sampling frequencies, the total duration of the filter could be covered with less than a thousand samples. In this context, considering a large filter with 1000 samples and high numerical precision (16 bits), around 16 filters can be stored in a 32-KB L1 cache memory beside each antenna. Considering modern memory designs, such cache memories have a very low access latency and would therefore not impact the performance of the protocol. For scaling to more filters per antenna, one could reduce the number of samples or numerical precision at the expense of a lower TR quality.
Moreover, TR filter calibration plays a vital role in error correction due to mismatches occurring in the channel response. Indeed, the real-time CIR could be deviated from the design phase due to discrepancies occurring due to chip manufacturing and packaging. Hence, feedback-based mechanisms can be employed for filter calibration to adjust the initial measurements, which is a common practice in testing the chips [12].